OpenStudy (anonymous):

Solve by substitution: 1/2x+y=9 3x-4y=-6

OpenStudy (zephyr141):

simplify either equation to x or y. then plug your equation into the corresponding variable for the other equation and solve for the only variable in that new equation. then put your results back into the other equation before you solved for the variable and solve for the unknown variable. i think this is the procedure to solve by substitution.

OpenStudy (anonymous):

um, okay?

OpenStudy (zephyr141):

ok take the first equation. solve for x in that one.

OpenStudy (anonymous):

Do i add or subtract the X's?

OpenStudy (zephyr141):

take the first equation \[\frac{ 1 }{ 2 }x+y=9\] and solve for x. meaning make it so that you only have x on one side of the equation like so x=....... to do this you must get x and the number attached to it alone by either adding/subtracting/multiplying/dividing other numbers or variables. in this case we see y alone so to get rid of it on the left side you have to subtract it. but remember since this is an equation it's saying that the left side is equal to the right side. so what you do to the left you have to do to the right to keep them equal. so you subtract y from the left side and the right side. it looks like this.\[\frac{ 1 }{ 2 }x+y(-y)=9(-y)\] which then becomes \[\frac{ 1 }{ x }x=9-y\]

OpenStudy (zephyr141):

crap. should be 1/2 not 1/x

OpenStudy (zephyr141):

now to find x in this equation you have to multiply both sides by 2 since you see the x getting divided by 2 in the left side. you should have \[x=18-2y\]

OpenStudy (zephyr141):

no take this and plug it into the second equation. \[3(here)-4y=-6\]

OpenStudy (zephyr141):

\[3(18-2y)-4y=-6\]which is \[54-6y-4y=-6\]

OpenStudy (zephyr141):

now combine like terms and divide both sides by the number you have on y.

OpenStudy (zephyr141):

this y value that you get will be the answer for both equations. now plug this y value into either \[3x-4y=-6\] or\[\frac{ 1 }{ 2 }x+y=9\] and solve for x to get your answer for x.